The Lions have won 40 of their first 74 baseball games. How many of the remaining 66 games must the team win to win 60% of their first games?
Answer by kunal_122(19)

ok let that x is the game needed to win
so 40+x/74=60%
that is 40+x/74=0.60
40+x=44.40
x=44.40-40
x=4.4
since u can play half game we will turnuate this to 4.
so they need 4 more games to win
is that what your textbook answer says?